OpenStudy (anonymous):

x²dy + √y²+4 dx + dy = 0

OpenStudy (primeralph):

Can I do this later? It's late.

OpenStudy (anonymous):

certainly, thanks

OpenStudy (anonymous):

elaborate what you meant

OpenStudy (loser66):

\[(x^2+1)dy+\sqrt{y^2+4}dx=0\\(x^2+1)dy=-\sqrt{y^2+4}dx\\\frac{dy}{\sqrt{y^2+4}}=\frac{dx}{x^2+1}\]Integrate both sides \[\int \frac{dy}{\sqrt{y^2+4}}=\int \frac{dx}{x^2+1}\] I don't think it's hard from this form.